# MOC-RVQ: Multilevel Codebook-assisted Digital Generative Semantic Communication

### v1

@@ Abstract
Vector quantization-based image semantic communication systems have successfully boosted transmission efficiency, but face a challenge with conflicting requirements between codebook design and digital constellation modulation. Traditional codebooks need a wide index range, while modulation favors few discrete states. To address this, we propose a multilevel generative semantic communication system with a two-stage training framework. In the first stage, we train a high-quality codebook, using a multi-head octonary codebook (MOC) to compress the index range. We also integrate a residual vector quantization (RVQ) mechanism for effective multilevel communication. In the second stage, a noise reduction block (NRB) based on Swin Transformer is introduced, coupled with the multilevel codebook from the first stage, serving as a high-quality semantic knowledge base (SKB) for generative feature restoration. Experimental results highlight MOC-RVQ’s superior performance over methods like BPG or JPEG, even without channel error correction coding.
@@
vector quantization, generative semantic communication, two-stage training, semantic knowledge base
@@ I Introduction
Traditional communication systems focus on the bit/symbol-level transmission, where the receiver seeks to minimize bit/symbol transmission errors to recover digital media [1]. This allows the design of communication system architecture to be separated from the vast and complex content of media. As a result, bit/symbol-based transmission has long been the prevailing paradigm in traditional communication systems.
Recently, propelled by the rapid development of deep learning technology, a new intelligent transmission paradigm known as semantic communication is gradually gaining prominence. In contrast to traditional transmission, semantic communication aims for the transmission of semantic fidelity [2]. For example, Dai et al. [3] integrate the nonlinear transform as a robust prior to efficiently extract source semantic features. Yang et al. [4] introduce a deep joint source-channel coding (JSCC) strategy designed for the transmission of images in wireless communication. Sun et al. [5] design a multi-level semantic coding and feature transmission mechanism powered by semantic knowledge base. However, these works suffer from efficiency issues, due to the high dimension of latent representation of raw data.
One solution to achieve a compact representation is vector quantization (VQ) technique [6] . Generally, VQ transforms semantic features into a series of indices, providing a more compact format that can be further converted into bits for transmission, then the receiver utilizes the received indices to reconstruct the semantic features through a pre-established learned codebook. Ultimately, the raw data or the task-oriented information can be decoded from these reconstructed semantic features. Nemati et al. [7] delve into the characteristics of VQ-VAE and adapt its training process to formulate a robust JSCC scheme against noisy wireless channels. Fu et al. [8] design a CNN-based transceiver to extract multi-scale semantic features and incorporate multi-scale semantic embedding spaces to facilitate feature quantization. To improve semantic representation in digital transmission under resource limitations, Guo et al. [9] introduce a novel non-linear quantization module with trainable levels for efficient feature extraction.
However, methods above suffer the following two issues:
  • Incompatible issue arises from the disparity between conventional vector quantization and digital constellation modulation. The former, aimed at achieving optimal image representation, employs a learnable codebook with a wide range of code indices. In contrast, the latter is more inclined to handle fewer states (e.g., 16-QAM, 64-QAM, etc.). While it is conceivable to regroup bits representing indices to align with modulation, such an adjustment also disrupts the explicit relationship between indices and constellation points. Note that this explicit relationship is crucial for preserving the local semantic relationship of the underlying neighbour code vectors, which directly affects the quality of reconstructed image.
@@
To tackle the incompatible issue, we introduce MOC-RVQ with two stage training framework. In Stage 1, we introduce (1) MOC, a variant of vector quantizer by constructing a multi-head octonary codebook (MOC) to compress discrete states to a limit of 8, allowing for a direct match with 64-QAM. (2) RVQ, a multi-level semantic transmission mechanism based on residual vector quantization (RVQ), which is helpful to compensate the negative effect of the quantization noise [10, 11]. These two components are integrated to form MOC-RVQ. Subsequently, in Stage 2, we propose a noise reduction block (NRB) and feature requantization to achieve generative feature restoration. For the mismatch issue, we adopt a heuristic codebook reordering algorithm, which is aimed at minimizing the semantic distance between underlying code vectors with nearby indices.
@@
  • We point out two inherent shortcomings of the existing VQ-based semantic communication system and propose a novel two-stage framework.
  • We introduce MOC-RVQ to align with the property of digital constellation modulation. Then noise reduction block (NRB) based on Swin Transformer and feature requantization are incorporated for generative semantic reconstruction.
  • We propose a codebook reordering algorithm, inspired by Gray code mapping, to alleviate the discrepancy in the local relationships among code vectors with nearby indices. This algorithm serves to enhance the overall robustness of the system against channel noise.
II System Model
In this section, we present a general VQ-based semantic communication framework, which we will subsequently extend to our two-stage framework in Section III. The core pipeline is depicted in Fig. 1. In this setup, both the transmitter and receiver employ a pre-established semantic codebook as semantic knowledge base (SKB) to achieve an efficient process for compressing and recovering of semantic features, ultimately leading to a reduction in communication overhead.
@@ II-A Transmitter
Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the semantic encoder $E$ (served as understanding knowledge base) is employed initially to transform $\mathbf{I}$ into a semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$. Subsequently, each grid feature $\mathbf{z}^{(i,j)}\in\mathbb{R}^{n_{q}}$ is substituted with the nearest entry in the learnable codebook $\mathcal{C}=\left\{\mathbf{e}_{k}\in\mathbb{R}^{n_{q}}\right\}_{k=0}^{N}$. This substitution yields the quantized feature $\mathbf{z}_{q}$ along with the corresponding code index sequence $\mathbf{s}\in\{0,1,\cdots,N-1\}^{h\cdot w}$, by the following equations:
@@
[…]
where $\mathbf{Q}(\cdot)$ is a quantization operation that maps the continuous encoded grid feature $\mathbf{z}^{(i,j)}$ into the quantized form $\mathbf{z}_{q}^{(i,j)}$ by looking up the closest codebook entry $\mathbf{e}_{k}$. Note that the code index sequence $\mathbf{s}$ can be served as a compact equivalent representation of the quantized feature $\mathbf{z}_{q}$. This implies that transmitting only the code index sequence $\mathbf{s}$ can effectively reduce communication overhead. Following decimal to binary transform and digital constellation modulation, $\mathbf{s}$ is converted as a transmitted complex signal $\mathbf{x}\in\mathbb{C}^{B\times 1}$, where $B$ represents the length of the transmitted signal.
@@
Figure 2: Left: The proposed two-stage framework. In Stage 1, we initially pretrain MOC-RVQ and other model components to generate compact representation of high-resolution image. Then the noise reduction block (NRB) is finetuned in Stage 2 to achieve feature restoration. Right: The architecture of the proposed MOC and MOC-RVQ. In MOC, each residual feature $\mathbf{r}_{d}$ is first divided into several heads, then the 8-state codebook is employed to quantize each head feature, and finally these features are concatenated together to form a single quantized feature. In MOC-RVQ, feature are recursive quantized by MOC to produce residual features and the corresponding code indices. Note that the output of MOC-RVQ is a summation over all quantized features $\mathbf{e}^{(d)}_{\mathbf{s}_{d}}$.
@@ II-C Receiver
After demodulation of the received signal $\mathbf{y}$, the receiver obtains $\hat{\mathbf{x}}$. Subsequently, the binary to decimal transform retrieves the indices $\mathbf{\hat{s}}$ from $\mathbf{\hat{x}}$ (Note that $\mathbf{\hat{s}}$ can be served as a prompt in a large language model, we utilize it to generate high quality reconstructed image). By selecting the corresponding embedding vectors from the codebook $\mathcal{C}$, the receiver reconstructs the semantic feature tensor $\mathbf{\hat{z}}_{q}$, we define this process as semantic feature retrieval $\mathbf{R}(\cdot)$. Finally, leveraging this retrieval semantic information, the semantic decoder $G$ (served as generative knowledge base) generates the reconstructed image $\mathbf{\hat{I}}$.
@@ III The Proposed Method
In this section, we introduce a novel two-stage framework to enhance vector quantization-based semantic communication. Subsequently, we delve into the description of codebook reordering, an algorithm employed to minimize the mismatch between local relationships of code indices and code vectors.
@@ III-A Stage 1: Pretraining of MOC-RVQ
The primary objective of Stage 1 is to train a high-quality encoder/decoder model and codebook without any channel noise, aiming for a compact representation of images. As depicted in Fig. 2 (Left), in Stage 1, an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is initially fed into the encoder $E$ to yield a continuous semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$. Subsequently, the MOC-RVQ layer quantizes this feature, resulting in the quantized feature $\mathbf{z}_{q}$ and an associated code indices sequence $\mathbf{s}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P\cdot D}$. Here, $P$ and $D$ represent the number of heads and the level of quantizations respectively, and $N$ is set to 8 to align with the few states prior of 64-QAM. Finally, the quantized feature $\mathbf{z}_{q}$ is fed into the decoder $G$ to reconstruct the image as $\mathbf{\hat{I}}$.
@@
Given a quantization level $d$, each residual feature $\mathbf{r}_{d}\in\mathbb{R}^{h\times h\times n_{q}}$ is initially divided into several heads $\mathbf{r}_{d,i}\in\mathbb{R}^{h\times h\times n_{q}/P}$ (typically set $P$ to 4 by default). Subsequently, an 8-state codebook is employed to quantize each head, yielding head-wise quantized features $\mathbf{e}_{d,i}$. These quantized features are then concatenated to form a single quantized feature $\mathbf{e}_{s_{d}}^{(d)}$. This distinctive codebook design for digital constellation modulation is termed as the multi-head octonary codebook (MOC). In contrast to conventional codebook with size $K$, our proposed MOC dramatically expands the feature matching space, generating $N^{P}$ discrete states through the combination of each head state.
@@
$\mathcal{MOC-RVQ}(\mathbf{z})=\left(\mathbf{s}_{1},\cdots,\mathbf{s}_{D}\right),$ (4)
where $\mathbf{s}_{d}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P}$ represents the code indices of $\mathbf{z}$ at level $d$, and $P$ denotes the number of heads. Given an initial residual $\mathbf{r}_{0}=\mathbf{z}$, the recursive process computes indices $s_{d}$ and the next residual $\mathbf{r}_{d}$ as follows:
@@
$\mathcal{L}^{\prime}_{VQ}(E,G,\mathcal{C})=\left\|\mathbf{\hat{I}}-\mathbf{I}\right\|_{1}+\sum_{d,h}\|\operatorname{sg}[\mathbf{r}_{d,h}]-\mathbf{e}_{d,h}\|_{2}^{2},$ (7)
where $sg[\cdot]$ represents the stop-gradient operation, and $\mathbf{r}_{d,h}$ and $\mathbf{e}_{d,h}$ denote the d-th quantization of the h-th head of residual and quantized features, respectively.
@@
where $CONV(\cdot)$ denotes a $1\times 1$ convolution layer to align the dimensions of $\mathbf{z}$, $\phi$ represents the pretrained VGG19 model, and $\gamma$ is set to 0.1. Additionally, in accordance with [12], perceptual loss and adversarial loss are employed for model pretraining.
@@ III-B Stage 2: Finetuning for Noise Reduction
Since Stage 1 is trained end-to-end without accounting for channel noise, leading to suboptimal reconstructed results in the presence of noise, we integrate a noise reduction block (NRB) based on a stack of residual Swin Transformer layers, as mentioned in [13] into decoder $G$ to conduct generative image reconstruction.
Building upon the pretrained model and codebook from Stage 1, as depicted in Fig. 2 (Left), Stage 2 introduces an physical communication channel module (refer to Section II-B for details) between transmitter and receiver for physical channel simulation. Then the noise reduction block is incorporated after the retrieval feature $\mathbf{\hat{z}}_{q}$ to refine semantic features, yielding the refined semantic feature $\mathbf{\hat{z}}_{r}$.
@@
To further enhance the quality of $\mathbf{\hat{z}}_{r}$, we propose requantizing it to obtain $\mathbf{\tilde{z}}_{r}$, formulated as follows:
@@
The intuition behind this lies in the fact that MOC is trained in the absence of any noise, establishing itself as a high-quality semantic knowledge base to inject high quality prior for better feature restoration.
@@
For NRB finetuning, we generate the noiseless quantized semantic feature of image $\mathbf{I}$ through $\mathbf{z}_{q}=\mathcal{MOC}(E(\mathbf{I}))$ as the ground-truth feature. This ground-truth feature is then used to calculate the Gram matrix loss and L2 loss for the refined semantic feature $\mathbf{\hat{z}}_{r}$ by the following formula:
@@
where $\psi(\cdot)$ is the flattened version of the Gram matrix of a given feature, and $\alpha$ is set to $0.25$. The first term of $\mathcal{L}_{NR}$, also known as style loss, has been demonstrated to be effective for improved feature recovery [13]. Note that the model and codebook trained in Stage 1 remain fixed, and we only finetune the NRB for noise reduction.

### v2

@@ Abstract
Vector quantization-based image semantic communication systems have successfully boosted transmission efficiency, but face challenges with conflicting requirements between codebook design and digital constellation modulation. Traditional codebooks need wide index ranges, while modulation favors few discrete states. To address this, we propose a multilevel generative semantic communication system with a two-stage training framework. In the first stage, we train a high-quality codebook, using a multi-head octonary codebook (MOC) to compress the index range. In addition, a residual vector quantization (RVQ) mechanism is also integrated for effective multilevel communication. In the second stage, a noise reduction block (NRB) based on Swin Transformer is introduced, coupled with the multilevel codebook from the first stage, serving as a high-quality semantic knowledge base (SKB) for generative feature restoration. Finally, to simulate modern image transmission scenarios, we employ a diverse collection of high-resolution 2K images as the test set. The experimental results consistently demonstrate the superior performance of MOC-RVQ over conventional methods such as BPG or JPEG. Additionally, MOC-RVQ achieves comparable performance to an analog JSCC scheme, while needing only one-sixth of the channel bandwidth ratio (CBR) and being directly compatible with digital transmission systems.
@@
Vector Quantization; Generative Semantic Communication; Semantic Knowledge Base; Image Transmission
@@ I Introduction
Propelled by the rapid development of deep learning technology, a new intelligent transmission paradigm known as semantic communication is gradually gaining prominence. In contrast to traditional transmission, semantic communication aims for the transmission of semantic fidelity [1]. For example, Dai et al. [2] integrate the nonlinear transform as a robust prior to efficiently extract source semantic features for image transmission. Sun et al. [3] design a multi-level semantic coding and feature transmission mechanism powered by semantic knowledge base. Wu et al. [4] propose channel denoising diffusion models (CDDM) for wireless communications by developing specific training and sampling algorithms tailored to the forward diffusion process and the reverse sampling process. However, these works suffer from efficiency issues, due to the high dimension of latent representation of raw data.
One solution to achieve compact representation is vector quantization (VQ) technique [5] . Generally, VQ transforms semantic features into a series of indices, providing a more compact format that can be further converted into bits for transmission, then the receiver utilizes the received indices to reconstruct the semantic features through a pre-trained semantic codebook. Ultimately, the raw data or the task-oriented information can be decoded from these reconstructed semantic features. For instance, Nemati et al. [6] delve into the characteristics of VQ-VAE and adapt its training process to formulate a robust JSCC scheme against noisy wireless channels. Fu et al. [7] design a CNN-based transceiver to extract multi-scale semantic features and incorporate multi-scale semantic embedding spaces to facilitate feature quantization.
However, the VQ methods mentioned above suffer the following two issues:
  • Incompatible issue arises from the disparity between conventional vector quantization and digital constellation modulation. The former, aimed at achieving optimal image representation, normally employs a learnable codebook with wide index ranges. In contrast, the latter is more inclined to handle fewer states (e.g., 16-QAM, 64-QAM, etc.). While it is conceivable to regroup bits representing indices to align with modulation, such an adjustment also disrupts the explicit relationship between indices and constellation points. Note that this explicit correspondence is crucial for preserving the local semantic relationship of the underlying neighbour code vectors, which directly affects the quality of reconstructed image.
@@
To tackle the incompatible issue, we introduce MOC-RVQ with a two-stage training framework. In Stage 1, we introduce: (1) MOC, a variant of vector quantizer by constructing a multi-head octonary codebook (MOC) to compress the range of indices to a limit of 8, allowing for a direct match with 64-QAM. (2) RVQ, a multi-level semantic transmission mechanism based on residual vector quantization (RVQ), which is helpful to compensate the negative effect of the quantization noise [8, 9]. We combine MOC and RVQ to form MOC-RVQ. In Stage 2, we introduce a noise reduction block (NRB) along with feature requantization to enable generative feature restoration. For the mismatch issue, we adopt a heuristic codebook reordering algorithm, which is aimed at minimizing the semantic distance between underlying code vectors with nearby indices.
@@
  • We point out two inherent shortcomings of the existing VQ-based semantic communication systems and propose a novel two-stage training framework.
  • We design MOC-RVQ to align with the preference of digital constellation modulation. Then the noise reduction block (NRB) based on Swin Transformer and feature requantization are incorporated for generative semantic reconstruction.
  • We develop a codebook reordering algorithm to mitigate inconsistencies in the local relationships among code vectors with adjacent indices. This algorithm serves to enhance the overall robustness of the system against channel noise.
II Preliminary
In this section, we present a general VQ-based semantic communication framework as preliminary. Note that this framework will be expanded and adjusted into our proposed two-stage framework in Section III. The main process is illustrated in Fig. 1. In this design, both the transmitter and receiver utilize a shared pre-trained semantic codebook to efficiently compress and reconstruct semantic features, thereby reducing communication overhead.
@@ II-A Transmitter
Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the semantic encoder $E$ (serves as understanding knowledge base) is employed initially to transform $\mathbf{I}$ into semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$. Subsequently, each grid feature $\mathbf{z}^{(i,j)}\in\mathbb{R}^{n_{q}}$ is substituted with the nearest vector in the learnable codebook $\mathcal{C}=\left\{\mathbf{e}_{k}\in\mathbb{R}^{n_{q}}\right\}_{k=0}^{N}$. This substitution yields the quantized feature $\mathbf{z}_{q}$ along with the corresponding code index sequence $\mathbf{s}\in\{0,1,\cdots,N-1\}^{h\cdot w}$, by the following equations:
[…]
@@
where, $\mathbf{Q}(\cdot)$ represents a quantization operation that maps the continuous encoded grid feature $\mathbf{z}^{(i,j)}$ to its quantized form $\mathbf{z}_{q}^{(i,j)}$ by selecting the closest codebook vector $\mathbf{e}_{k}$. Notably, the code index sequence $\mathbf{s}$ can serve as a compact equivalent representation of the quantized feature $\mathbf{z}_{q}$, meaning that transmitting only the code index sequence $\mathbf{s}$ can effectively reduce communication overhead. After applying decimal-to-binary transformation and digital constellation modulation, $\mathbf{s}$ is transformed into a transmitted complex signal $\mathbf{x}\in\mathbb{C}^{B\times 1}$, where $B$ denotes the length of the transmitted signal.
@@
Figure 2: Left: The proposed two-stage framework. In Stage 1, we initially pretrain MOC-RVQ and other model components to generate compact representation of high-resolution image. Then the noise reduction block (NRB) is finetuned in Stage 2 to achieve feature restoration. Right: The architecture of the proposed MOC and MOC-RVQ. In MOC, each residual feature $\mathbf{r}_{d-1}$ is first divided into several heads, then the 8-state codebooks are employed to quantize each head feature, and finally these features are concatenated together to form a single quantized feature. In MOC-RVQ, feature are recursive quantized by MOCs to produce residual features and the corresponding code indices. Note that the output of MOC-RVQ is a summation over all quantized features $\mathbf{e}^{(d)}_{\mathbf{s}_{d}}$.
@@ II-C Receiver
Following the demodulation of the received signal $\mathbf{y}$ and the subsequent binary-to-decimal transformation, the receiver retrieves the indices $\mathbf{\hat{s}}$ from $\mathbf{y}$. It is noteworthy that $\mathbf{\hat{s}}$ can be analogized to the prompt in large language models, and we employ it to generate high-quality reconstructed images. Specifically, by selecting the corresponding embedding vectors from the codebook $\mathcal{C}$ using the received indices $\mathbf{\hat{s}}$, the receiver reconstructs the semantic feature $\mathbf{\hat{z}}_{q}$, we define this process as semantic feature retrieval $\mathbf{R}(\cdot)$. Then, leveraging the retrieval semantic feature $\mathbf{\hat{z}}_{q}$, the semantic decoder $G$ (serves as generative knowledge base) generates the reconstructed image $\mathbf{\hat{I}}$.
@@ III The Proposed Method
In this section, we expand the pipeline mentioned in Section II and introduce a novel two-stage training framework to enhance VQ-based semantic communication. Subsequently, we delve into the description of codebook reordering, an algorithm employed to minimize the mismatch between local relationships of code indices and code vectors.
@@ III-A Stage 1: Pretraining of MOC-RVQ
The primary objective of Stage 1 is to train a high-quality encoder/decoder model and multilevel codebook without any channel noise, aiming for a compact representation of image. As depicted in Fig. 2 (Left), in Stage 1, an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is initially fed into the encoder $E$ to yield a continuous semantic feature $\mathbf{z}\in\mathbb{R}^{h\times w\times n_{q}}$. Subsequently, the MOC-RVQ layer quantizes this feature, resulting in the quantized feature $\mathbf{z}_{q}$ and the associated code indices sequence $\mathbf{s}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P\cdot D}$. Here, $P$ and $D$ represent the number of heads and the level of quantizations respectively, and $N$ is set to 8 to directly align with the few states preference of 64-QAM. Finally, the quantized feature $\mathbf{z}_{q}$ is fed into the decoder $G$ to reconstruct the image as $\mathbf{\hat{I}}$.
@@
Given a quantization level $d$, each residual feature $\mathbf{r}_{d-1}\in\mathbb{R}^{h\times w\times n_{q}}$ is initially divided into several heads $\mathbf{r}_{d-1,i}\in\mathbb{R}^{h\times w\times n_{q}/P}$ (typically set $P$ to 4 by default). Subsequently, the 8-state codebooks are employed to quantize each head separately, resulting in head-wise quantized features $\mathbf{e}_{d,i}$. These quantized features are then concatenated to form a single quantized feature $\mathbf{e}_{s_{d}}^{(d)}$. We name this distinctive codebook design for aliging the preference of digital constellation modulation as multi-head octonary codebook (MOC). It is worth noting that, unlike a conventional codebook of size $K$, our proposed MOC significantly expands the feature matching space by generating $N^{P}$ discrete states through the combination of individual head states.
@@
$\mathcal{MOC-RVQ}(\mathbf{z})=\left(\mathbf{s}_{1},\cdots,\mathbf{s}_{d},\cdots,\mathbf{s}_{D}\right),$ (4)
where $\mathbf{s}_{d}\in\{0,\cdots,N-1\}^{h\cdot w\cdot P}$ represents the code indices of $\mathbf{z}$ at level $d$, and $P$ denotes the number of heads. More specifically, given an initial residual $\mathbf{r}_{0}=\mathbf{z}$, the recursive process computes indices $s_{d}$ and the next residual $\mathbf{r}_{d}$ as follows:
@@
$\mathcal{L}^{\prime}_{VQ}(E,G,\mathcal{C})=\left\|\mathbf{\hat{I}}-\mathbf{I}\right\|_{1}+\sum_{d,h}\|\operatorname{sg}[\mathbf{r}_{d-1,h}]-\mathbf{e}_{d,h}\|_{2}^{2},$ (7)
where $sg[\cdot]$ represents the stop-gradient operation. Note that the first term of $\mathcal{L}^{\prime}_{VQ}$ stands for the image-level $L_{1}$ loss, while the second term is employed for updating the proposed multilevel octonary codebooks.
@@
where $CONV(\cdot)$ denotes a $1\times 1$ convolution layer to align the dimensions of $\mathbf{z}$, $\phi$ represents the pretrained VGG19 model, and $\gamma$ is set to 0.1. Additionally, in accordance with [10], perceptual loss and adversarial loss are also employed for model pretraining.
@@ III-B Stage 2: Finetuning for Noise Reduction
Since Stage 1 is trained end-to-end without considering any channel noise, leading to suboptimal reconstructed results in the presence of noise, we integrate a noise reduction block (NRB) based on a stack of residual Swin Transformer layers, as mentioned in [11] into decoder $G$ to conduct generative image reconstruction.
Building upon the pretrained model and multilevel codebook from Stage 1, as depicted in Fig. 2 (Left), Stage 2 introduces a physical communication channel (refer to Section II-B for details) between transmitter and receiver for physical channel simulation. Then the proposed NRB is incorporated after the retrieval noisy feature $\mathbf{\hat{z}}_{q}$ to refine semantic features, yielding the refined semantic feature $\mathbf{\hat{z}}_{r}$.
@@
To further enhance the quality of $\mathbf{\hat{z}}_{r}$, we propose requantizing it to obtain $\mathbf{\tilde{z}}_{r}$ by:
@@
The intuition behind this lies in the fact that MOC-RVQ is trained in the absence of any noise, establishing itself as a high-quality semantic knowledge base to inject high quality prior for better feature restoration.
@@
For NRB finetuning, we first generate the noiseless quantized semantic feature of image $\mathbf{I}$ through $\mathbf{z}_{q}=\mathcal{MOC-RVQ}(E(\mathbf{I}))$ as ground-truth feature. Then, this ground-truth feature is utilized to calculate the Gram matrix loss and L2 loss for the refined semantic feature $\mathbf{\hat{z}}_{r}$ by:
@@
where $\psi(\cdot)$ is the flattened version of the Gram matrix of a given feature, and $\alpha$ is set to $0.25$. The first term of $\mathcal{L}_{NR}$, also known as style loss, has been demonstrated to be effective for improved feature recovery [11]. Moreover, we incorporate image-level $L_{1}$ loss for training supervision. It is also noteworthy that the autoencoder and codebook trained in Stage 1 remain fixed, and we only finetune the NRB for noise reduction.